OpenStudy (anonymous):

y=3x+9 Y=2x+6 how do i do this

OpenStudy (danjs):

y = both those things, you can maybe just compare y=y,. or 3x+9 = 2x + 6

OpenStudy (danjs):

or , take (-1) times one of those equations, and then add them together y + (-y) goes to zero for the left side

OpenStudy (pawanyadav):

Subtract eq 2 from eq 1, LHS from LHS and RHS from RHS, We get, y-y=3x+9-2x-6 0=x+3 Subtract 3 both side 0-3=x+3-3 -3=x So x=-3